# Visual Writing: Writing by Manipulating Visual Representations of Stories

### v1

@@ Acknowledgements.
- Dang et al. (2022a) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek. 2022a. Beyond Text Generation: Supporting Writers with Continuous Automatic Text Summaries. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. ACM, Bend OR USA, 1–13. https://doi.org/10.1145/3526113.3545672
- Dang et al. (2023) Hai Dang, Sven Goller, Florian Lehmann, and Daniel Buschek. 2023. Choice Over Control: How Users Write with Large Language Models Using Diegetic and Non-Diegetic Prompting. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–17. https://doi.org/10.1145/3544548.3580969 arXiv:2303.03199 [cs]
- Dang et al. (2022b) Hai Dang, Lukas Mecke, Florian Lehmann, Sven Goller, and Daniel Buschek. 2022b. How to Prompt? Opportunities and Challenges of Zero- and Few-Shot Learning for Human-AI Interaction in Creative Applications of Generative Models. https://arxiv.org/abs/2209.01390v1.
- Freytag (1895) Gustav Freytag. 1895. Technique of the drama: An exposition of dramatic composition and art. S. Griggs.
- Gobert and Beaudouin-Lafon (2023) Camille Gobert and Michel Beaudouin-Lafon. 2023. Lorgnette: Creating Malleable Code Projections. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3586183.3606817
- Hempel et al. (2019) Brian Hempel, Justin Lubin, and Ravi Chugh. 2019. Sketch-n-Sketch: Output-Directed Programming for SVG. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 281–292. https://doi.org/10.1145/3332165.3347925
@@
- Hoque et al. (2023) Md Naimul Hoque, Bhavya Ghai, Kari Kraus, and Niklas Elmqvist. 2023. Portrayal: Leveraging NLP and Visualization for Analyzing Fictional Characters. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. ACM, Pittsburgh PA USA, 74–94. https://doi.org/10.1145/3563657.3596000
@@
- Hutchins et al. (1985) Edwin L. Hutchins, James D. Hollan, and Donald A. Norman. 1985. Direct Manipulation Interfaces. Human–Computer Interaction 1, 4 (Dec. 1985), 311–338. https://doi.org/10.1207/s15327051hci0104_2
@@
- Kempen et al. (1987) Gerard Kempen, G. Ambeek, Peter Desain, L. Konst, and Koenraad De Smedt. 1987. Author Environments: Fifth Generation Text Processors. (Jan. 1987).
@@
- Kim et al. (2018) Nam Wook Kim, Benjamin Bach, Hyejin Im, Sasha Schriber, Markus Gross, and Hanspeter Pfister. 2018. Visualizing Nonlinear Narratives with Story Curves. IEEE Transactions on Visualization and Computer Graphics 24, 1 (Jan. 2018), 595–604. https://doi.org/10.1109/TVCG.2017.2744118
@@
- Ko and Myers (2006) Amy J. Ko and Brad A. Myers. 2006. Barista: An Implementation Framework for Enabling New Tools, Interaction Techniques and Views in Code Editors. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’06). Association for Computing Machinery, New York, NY, USA, 387–396. https://doi.org/10.1145/1124772.1124831
@@
- Lee et al. (2024) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A. Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L. C. Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Agnia Sergeyuk, Antonette Shibani, Disha Shrivastava, Lila Shroff, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia H. Rho, Shannon Zejiang Shen, and Pao Siangliulue. 2024. A Design Space for Intelligent and Interactive Writing Assistants. https://doi.org/10.1145/3613904.3642697 arXiv:2403.14117 [cs]
@@
- Maharjan et al. (2018) Suraj Maharjan, Sudipta Kar, Manuel Montes, Fabio A. González, and Thamar Solorio. 2018. Letting Emotions Flow: Success Prediction by Modeling the Flow of Emotions in Books. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 259–265. https://doi.org/10.18653/v1/N18-2042
- Marshall (2014) Colin Marshall. 2014. How J.K. Rowling Plotted Harry Potter with a Hand-Drawn Spreadsheet | Open Culture.
@@
- Marti et al. (2018) Marcel Marti, Jodok Vieli, Wojciech Witoń, Rushit Sanghrajka, Daniel Inversini, Diana Wotruba, Isabel Simo, Sasha Schriber, Mubbasir Kapadia, and Markus Gross. 2018. CARDINAL: Computer Assisted Authoring of Movie Scripts. In Proceedings of the 23rd International Conference on Intelligent User Interfaces (IUI ’18). Association for Computing Machinery, New York, NY, USA, 509–519. https://doi.org/10.1145/3172944.3172972
@@
- Masson et al. (2023) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2023. Statslator: Interactive Translation of NHST and Estimation Statistics Reporting Styles in Scientific Documents. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–14. https://doi.org/10.1145/3586183.3606762
@@
- Masson et al. (2024) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2024. DirectGPT: A Direct Manipulation Interface to Interact with Large Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3613904.3642462
@@
- Masson et al. (2022) Damien Masson, Jo Vermeulen, George Fitzmaurice, and Justin Matejka. 2022. Supercharging Trial-and-Error for Learning Complex Software Applications. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3491102.3501895
@@
- Mayer et al. (2018) Mikaël Mayer, Viktor Kuncak, and Ravi Chugh. 2018. Bidirectional Evaluation with Direct Manipulation. Implentation for: Bidirectional Evaluation with Direct Manipulation 2, OOPSLA (Oct. 2018), 127:1–127:28. https://doi.org/10.1145/3276497
- Meyer et al. (2023) Jesse G. Meyer, Ryan J. Urbanowicz, Patrick C. N. Martin, Karen O’Connor, Ruowang Li, Pei-Chen Peng, Tiffani J. Bright, Nicholas Tatonetti, Kyoung Jae Won, Graciela Gonzalez-Hernandez, and Jason H. Moore. 2023. ChatGPT and Large Language Models in Academia: Opportunities and Challenges. BioData Mining 16, 1 (July 2023), 20. https://doi.org/10.1186/s13040-023-00339-9
- Morris et al. (2024) Meredith Ringel Morris, Jascha Sohl-dickstein, Noah Fiedel, Tris Warkentin, Allan Dafoe, Aleksandra Faust, Clement Farabet, and Shane Legg. 2024. Levels of AGI for Operationalizing Progress on the Path to AGI. https://doi.org/10.48550/arXiv.2311.02462 arXiv:2311.02462 [cs]
@@
- Myers (1986) B. A. Myers. 1986. Visual Programming, Programming by Example, and Program Visualization: A Taxonomy. SIGCHI Bull. 17, 4 (April 1986), 59–66. https://doi.org/10.1145/22339.22349
- Neuwirth and Kaufer (1989) C. M. Neuwirth and D. S. Kaufer. 1989. The Role of External Representation in the Writing Process: Implications for the Design of Hypertext-Based Writing Tools. In Proceedings of the Second Annual ACM Conference on Hypertext - HYPERTEXT ’89. ACM Press, Pittsburgh, Pennsylvania, United States, 319–341. https://doi.org/10.1145/74224.74250
@@
- Paivio (1990) Allan Paivio. 1990. Mental Representations: A Dual Coding Approach. Oxford University Press.
- Pope (1998) Thomas Pope. 1998. Good scripts, bad scripts: learning the craft of screenwriting through 25 of the best and worst films in history. Three Rivers Press.
- Qiang and Bingjie (2016) Lu Qiang and Chai Bingjie. 2016. StoryCake: A Hierarchical Plot Visualization Method for Storytelling in Polar Coordinates. In 2016 International Conference on Cyberworlds (CW). IEEE, Chongqing, China, 211–218. https://doi.org/10.1109/CW.2016.43
@@
- Reagan et al. (2016) Andrew J. Reagan, Lewis Mitchell, Dilan Kiley, Christopher M. Danforth, and Peter Sheridan Dodds. 2016. The Emotional Arcs of Stories Are Dominated by Six Basic Shapes. EPJ Data Science 5, 1 (Dec. 2016), 1–12. https://doi.org/10.1140/epjds/s13688-016-0093-1
- Reza et al. (2024) Mohi Reza, Nathan Laundry, Ilya Musabirov, Peter Dushniku, Zhi Yuan ”Michael” Yu, Kashish Mittal, Tovi Grossman, Michael Liut, Anastasia Kuzminykh, and Joseph Jay Williams. 2024. ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks Using Large Language Models. https://doi.org/10.1145/3613904.3641899 arXiv:2310.00117 [cs]
@@
- Shneiderman (1983) Shneiderman. 1983. Direct Manipulation: A Step Beyond Programming Languages. Computer 16, 8 (Aug. 1983), 57–69. https://doi.org/10.1109/MC.1983.1654471
- Shneiderman (1982) Ben Shneiderman. 1982. The Future of Interactive Systems and the Emergence of Direct Manipulation†. Behaviour & Information Technology 1, 3 (July 1982), 237–256. https://doi.org/10.1080/01449298208914450
- Shneiderman et al. (2006) Ben Shneiderman, Gerhard Fischer, Mary Czerwinski, Mitch Resnick, Brad Myers, Linda Candy, Ernest Edmonds, Mike Eisenberg, Elisa Giaccardi, Tom Hewett, Pamela Jennings, Bill Kules, Kumiyo Nakakoji, Jay Nunamaker, Randy Pausch, Ted Selker, Elisabeth Sylvan, and Michael Terry. 2006. Creativity Support Tools: Report From a U.S. National Science Foundation Sponsored Workshop. International Journal of Human–Computer Interaction 20, 2 (May 2006), 61–77. https://doi.org/10.1207/s15327590ijhc2002_1
@@
- Suh et al. (2024) Sangho Suh, Meng Chen, Bryan Min, Toby Jia-Jun Li, and Haijun Xia. 2024. Luminate: Structured Generation and Exploration of Design Space with Large Language Models for Human-AI Co-Creation. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 644, 26 pages. https://doi.org/10.1145/3613904.3642400
@@
- Tanahashi and Kwan-Liu Ma (2012) Y. Tanahashi and Kwan-Liu Ma. 2012. Design Considerations for Optimizing Storyline Visualizations. IEEE Transactions on Visualization and Computer Graphics 18, 12 (Dec. 2012), 2679–2688. https://doi.org/10.1109/TVCG.2012.212
@@
- Watson et al. (2019) Katie Watson, Samuel S. Sohn, Sasha Schriber, Markus Gross, Carlos Manuel Muniz, and Mubbasir Kapadia. 2019. StoryPrint: An Interactive Visualization of Stories. In Proceedings of the 24th International Conference on Intelligent User Interfaces (IUI ’19). Association for Computing Machinery, New York, NY, USA, 303–311. https://doi.org/10.1145/3301275.3302302
- Weidinger et al. (2021) Laura Weidinger, John Mellor, Maribeth Rauh, Conor Griffin, Jonathan Uesato, Po-Sen Huang, Myra Cheng, Mia Glaese, Borja Balle, Atoosa Kasirzadeh, Zac Kenton, Sasha Brown, Will Hawkins, Tom Stepleton, Courtney Biles, Abeba Birhane, Julia Haas, Laura Rimell, Lisa Anne Hendricks, William Isaac, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2021. Ethical and Social Risks of Harm from Language Models. https://doi.org/10.48550/arXiv.2112.04359 arXiv:2112.04359 [cs]
- Williams ([n. d.]) Eric Williams. [n. d.]. How to View and Appreciate Great Movies. https://www.thegreatcourses.com/courses/how-to-view-and-appreciate-great-movies
@@
- Yuan et al. (2022) Ann Yuan, Andy Coenen, Emily Reif, and Daphne Ippolito. 2022. Wordcraft: Story Writing With Large Language Models. In Proceedings of the 27th International Conference on Intelligent User Interfaces (IUI ’22). Association for Computing Machinery, New York, NY, USA, 841–852. https://doi.org/10.1145/3490099.3511105
@@
- Zhang et al. (2023) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023. VISAR: A Human-AI Argumentative Writing Assistant with Visual Programming and Rapid Draft Prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco CA USA, 1–30. https://doi.org/10.1145/3586183.3606800
@@
- Zhu (2013) Xiaojin Zhu. 2013. Persistent Homology: An Introduction and a New Text Representation for Natural Language Processing. In IJCAI International Joint Conference on Artificial Intelligence. 1953–1959.

### v2

@@ Acknowledgements.
- Dang et al. (2022) Hai Dang, Karim Benharrak, Florian Lehmann, and Daniel Buschek. 2022. Beyond Text Generation: Supporting Writers with Continuous Automatic Text Summaries. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. ACM, Bend OR USA, 1–13. doi:10.1145/3526113.3545672
- Dang et al. (2023) Hai Dang, Sven Goller, Florian Lehmann, and Daniel Buschek. 2023. Choice Over Control: How Users Write with Large Language Models Using Diegetic and Non-Diegetic Prompting. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–17. doi:10.1145/3544548.3580969
- Flower and Hayes (1981) Linda Flower and John R. Hayes. 1981. A Cognitive Process Theory of Writing. College Composition and Communication 32, 4 (1981), 365–387. arXiv:356600 doi:10.2307/356600
- Freytag (2004) Gustav Freytag. 2004. Technique of the Drama: An Exposition of Dramatic Composition and Art. University Press of the Pacific.
- Genette et al. (1990) Gérard Genette, Nitsa Ben-Ari, and Brian McHale. 1990. Fictional narrative, factual narrative. Poetics today 11, 4 (1990), 755–774.
- Gobert and Beaudouin-Lafon (2023) Camille Gobert and Michel Beaudouin-Lafon. 2023. Lorgnette: Creating Malleable Code Projections. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–16. doi:10.1145/3586183.3606817
- Hempel et al. (2019) Brian Hempel, Justin Lubin, and Ravi Chugh. 2019. Sketch-n-Sketch: Output-Directed Programming for SVG. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (UIST ’19). Association for Computing Machinery, New York, NY, USA, 281–292. doi:10.1145/3332165.3347925
@@
- Hoque et al. (2023) Md Naimul Hoque, Bhavya Ghai, Kari Kraus, and Niklas Elmqvist. 2023. Portrayal: Leveraging NLP and Visualization for Analyzing Fictional Characters. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. ACM, Pittsburgh PA USA, 74–94. doi:10.1145/3563657.3596000
- Hsu et al. (2019a) Ting-Yao Hsu, Yen-Chia Hsu, and Ting-Hao (Kenneth) Huang. 2019a. On How Users Edit Computer-Generated Visual Stories. In Extended Abstracts of the 2019 CHI Conference on Human Factors in Computing Systems (CHI EA ’19). Association for Computing Machinery, New York, NY, USA, 1–6. doi:10.1145/3290607.3312965
- Hsu et al. (2019b) Ting-Yao Hsu, Chieh-Yang Huang, Yen-Chia Hsu, and Ting-Hao Huang. 2019b. Visual Story Post-Editing. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Anna Korhonen, David Traum, and Lluís Màrquez (Eds.). Association for Computational Linguistics, Florence, Italy, 6581–6586. doi:10.18653/v1/P19-1658
- Huang et al. (2016) Ting-Hao Kenneth Huang, Francis Ferraro, Nasrin Mostafazadeh, Ishan Misra, Aishwarya Agrawal, Jacob Devlin, Ross Girshick, Xiaodong He, Pushmeet Kohli, Dhruv Batra, C. Lawrence Zitnick, Devi Parikh, Lucy Vanderwende, Michel Galley, and Margaret Mitchell. 2016. Visual Storytelling. In Proceedings of the 2016 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kevin Knight, Ani Nenkova, and Owen Rambow (Eds.). Association for Computational Linguistics, San Diego, California, 1233–1239. doi:10.18653/v1/N16-1147
- Hulstein et al. (2023) Golina Hulstein, Vanessa Peña-Araya, and Anastasia Bezerianos. 2023. Geo-Storylines: Integrating Maps into Storyline Visualizations. IEEE Transactions on Visualization and Computer Graphics 29, 1 (Jan. 2023), 994–1004. doi:10.1109/TVCG.2022.3209480
- Hume (2014) Kathryn Hume. 2014. Fantasy and Mimesis: Responses to Reality in Western Literature. Routledge, Abingdon New York.
- Huot et al. (2025) Fantine Huot, Reinald Kim Amplayo, Jennimaria Palomaki, Alice Shoshana Jakobovits, Elizabeth Clark, and Mirella Lapata. 2025. Agents’ Room: Narrative Generation through Multi-step Collaboration. arXiv:2410.02603 [cs] doi:10.48550/arXiv.2410.02603
@@
- Hutchins et al. (1985) Edwin L. Hutchins, James D. Hollan, and Donald A. Norman. 1985. Direct Manipulation Interfaces. Human–Computer Interaction 1, 4 (Dec. 1985), 311–338. doi:10.1207/s15327051hci0104_2
@@
- Kempen et al. (1987) Gerard Kempen, Gert Anbeek, Peter Desain, Leo Konst, and Koenraad De Smedt. 1987. Author Environments: Fifth Generation Text Processors. In ESPRIT’86: Results and Achievements. Elsevier Science Publishers, Brussels, Belgium, 8.
@@
- Kim et al. (2018) Nam Wook Kim, Benjamin Bach, Hyejin Im, Sasha Schriber, Markus Gross, and Hanspeter Pfister. 2018. Visualizing Nonlinear Narratives with Story Curves. IEEE Transactions on Visualization and Computer Graphics 24, 1 (Jan. 2018), 595–604. doi:10.1109/TVCG.2017.2744118
@@
- Ko and Myers (2006) Amy J. Ko and Brad A. Myers. 2006. Barista: An Implementation Framework for Enabling New Tools, Interaction Techniques and Views in Code Editors. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’06). Association for Computing Machinery, New York, NY, USA, 387–396. doi:10.1145/1124772.1124831
@@
- Lee et al. (2024) Mina Lee, Katy Ilonka Gero, John Joon Young Chung, Simon Buckingham Shum, Vipul Raheja, Hua Shen, Subhashini Venugopalan, Thiemo Wambsganss, David Zhou, Emad A. Alghamdi, Tal August, Avinash Bhat, Madiha Zahrah Choksi, Senjuti Dutta, Jin L. C. Guo, Md Naimul Hoque, Yewon Kim, Simon Knight, Seyed Parsa Neshaei, Agnia Sergeyuk, Antonette Shibani, Disha Shrivastava, Lila Shroff, Jessi Stark, Sarah Sterman, Sitong Wang, Antoine Bosselut, Daniel Buschek, Joseph Chee Chang, Sherol Chen, Max Kreminski, Joonsuk Park, Roy Pea, Eugenia H. Rho, Shannon Zejiang Shen, and Pao Siangliulue. 2024. A Design Space for Intelligent and Interactive Writing Assistants. arXiv:2403.14117 [cs] doi:10.1145/3613904.3642697
- Liveley (2021) Genevieve Liveley. 2021. Narratology. Oxford University Press, Oxford.
- Madden (1988) David Madden. 1988. Revising Fiction: A Handbook for Writers. Plume, New York.
@@
- Maharjan et al. (2018) Suraj Maharjan, Sudipta Kar, Manuel Montes, Fabio A. González, and Thamar Solorio. 2018. Letting Emotions Flow: Success Prediction by Modeling the Flow of Emotions in Books. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 259–265. doi:10.18653/v1/N18-2042
@@
- Marti et al. (2018) Marcel Marti, Jodok Vieli, Wojciech Witoń, Rushit Sanghrajka, Daniel Inversini, Diana Wotruba, Isabel Simo, Sasha Schriber, Mubbasir Kapadia, and Markus Gross. 2018. CARDINAL: Computer Assisted Authoring of Movie Scripts. In Proceedings of the 23rd International Conference on Intelligent User Interfaces (IUI ’18). Association for Computing Machinery, New York, NY, USA, 509–519. doi:10.1145/3172944.3172972
- Masson et al. (2025) Damien Masson, Young-Ho Kim, and Fanny Chevalier. 2025. Textoshop: Interactions Inspired by Drawing Software to Facilitate Text Editing. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 775, 14 pages. doi:10.1145/3706598.3713862
@@
- Masson et al. (2023) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2023. Statslator: Interactive Translation of NHST and Estimation Statistics Reporting Styles in Scientific Documents. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3586183.3606762
@@
- Masson et al. (2024) Damien Masson, Sylvain Malacria, Géry Casiez, and Daniel Vogel. 2024. DirectGPT: A Direct Manipulation Interface to Interact with Large Language Models. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–16. doi:10.1145/3613904.3642462
@@
- Masson et al. (2022) Damien Masson, Jo Vermeulen, George Fitzmaurice, and Justin Matejka. 2022. Supercharging Trial-and-Error for Learning Complex Software Applications. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–13. doi:10.1145/3491102.3501895
@@
- Mayer et al. (2018) Mikaël Mayer, Viktor Kuncak, and Ravi Chugh. 2018. Bidirectional Evaluation with Direct Manipulation. Implentation for: Bidirectional Evaluation with Direct Manipulation 2, OOPSLA (Oct. 2018), 127:1–127:28. doi:10.1145/3276497
@@
- Myers (1986) B. A. Myers. 1986. Visual Programming, Programming by Example, and Program Visualization: A Taxonomy. SIGCHI Bull. 17, 4 (April 1986), 59–66. doi:10.1145/22339.22349
- Neuwirth and Kaufer (1989) C. M. Neuwirth and D. S. Kaufer. 1989. The Role of External Representation in the Writing Process: Implications for the Design of Hypertext-Based Writing Tools. In Proceedings of the Second Annual ACM Conference on Hypertext - HYPERTEXT ’89. ACM Press, Pittsburgh, Pennsylvania, United States, 319–341. doi:10.1145/74224.74250
@@
- North and Shneiderman (2001) Chris North and Ben Shneiderman. 2001. A Taxonomy of Multiple Window Coordinations. (2001), 8.
- Norton (2011) Scott Norton. 2011. Developmental Editing: A Handbook for Freelancers, Authors, and Publishers (illustrated edition ed.). University of Chicago Press, Chicago London.
@@
- Paivio (1990) Allan Paivio. 1990. Mental Representations: A Dual Coding Approach. Oxford University Press, Oxford.
- Pope (1998) Tom Pope. 1998. Good Scripts, Bad Scripts: Learning the Craft of Screenwriting Through 25 of the Best and Worst Films in Hi Story. Three Rivers Press, New York.
- Propp and Pirkova-Jakobson (2010) V. Propp and Svatava Pirkova-Jakobson. 2010. Morphology of the Folk Tale. University of Texas Press.
- Qiang and Bingjie (2016) Lu Qiang and Chai Bingjie. 2016. StoryCake: A Hierarchical Plot Visualization Method for Storytelling in Polar Coordinates. In 2016 International Conference on Cyberworlds (CW). IEEE, Chongqing, China, 211–218. doi:10.1109/CW.2016.43
@@
- Reagan et al. (2016) Andrew J. Reagan, Lewis Mitchell, Dilan Kiley, Christopher M. Danforth, and Peter Sheridan Dodds. 2016. The Emotional Arcs of Stories Are Dominated by Six Basic Shapes. EPJ Data Science 5, 1 (Dec. 2016), 1–12. doi:10.1140/epjds/s13688-016-0093-1
- Reza et al. (2024) Mohi Reza, Nathan Laundry, Ilya Musabirov, Peter Dushniku, Zhi Yuan ”Michael” Yu, Kashish Mittal, Tovi Grossman, Michael Liut, Anastasia Kuzminykh, and Joseph Jay Williams. 2024. ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks Using Large Language Models. arXiv:2310.00117 [cs] doi:10.1145/3613904.3641899
@@
- Shneiderman (1983) Shneiderman. 1983. Direct Manipulation: A Step Beyond Programming Languages. Computer 16, 8 (Aug. 1983), 57–69. doi:10.1109/MC.1983.1654471
- Shneiderman (1982) Ben Shneiderman. 1982. The Future of Interactive Systems and the Emergence of Direct Manipulation†. Behaviour & Information Technology 1, 3 (July 1982), 237–256. doi:10.1080/01449298208914450
- Shneiderman et al. (2006) Ben Shneiderman, Gerhard Fischer, Mary Czerwinski, Mitch Resnick, Brad Myers, Linda Candy, Ernest Edmonds, Mike Eisenberg, Elisa Giaccardi, Tom Hewett, Pamela Jennings, Bill Kules, Kumiyo Nakakoji, Jay Nunamaker, Randy Pausch, Ted Selker, Elisabeth Sylvan, and Michael Terry. 2006. Creativity Support Tools: Report From a U.S. National Science Foundation Sponsored Workshop. International Journal of Human–Computer Interaction 20, 2 (May 2006), 61–77. doi:10.1207/s15327590ijhc2002_1
@@
- Suh et al. (2024) Sangho Suh, Meng Chen, Bryan Min, Toby Jia-Jun Li, and Haijun Xia. 2024. Luminate: Structured Generation and Exploration of Design Space with Large Language Models for Human-AI Co-Creation. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 644, 26 pages. doi:10.1145/3613904.3642400
@@
- Tanahashi and Kwan-Liu Ma (2012) Y. Tanahashi and Kwan-Liu Ma. 2012. Design Considerations for Optimizing Storyline Visualizations. IEEE Transactions on Visualization and Computer Graphics 18, 12 (Dec. 2012), 2679–2688. doi:10.1109/TVCG.2012.212
- Todorov (1971) Tzvetan Todorov. 1971. The 2 Principles of Narrative. Diacritics 1, 1 (1971), 37–44. arXiv:464558 doi:10.2307/464558
- Wang et al. (2018) Jing Wang, Jianlong Fu, Jinhui Tang, Zechao Li, and Tao Mei. 2018. Show, Reward and Tell: Automatic Generation of Narrative Paragraph from Photo Stream by Adversarial Training. In Proceedings of the Thirty-Second AAAI Conference on Artificial Intelligence and Thirtieth Innovative Applications of Artificial Intelligence Conference and Eighth AAAI Symposium on Educational Advances in Artificial Intelligence (AAAI’18/IAAI’18/EAAI’18). AAAI Press, New Orleans, Louisiana, USA, 7396–7403.
- Wang Baldonado et al. (2000) Michelle Q. Wang Baldonado, Allison Woodruff, and Allan Kuchinsky. 2000. Guidelines for using multiple views in information visualization. In Proceedings of the working conference on Advanced visual interfaces. ACM, Palermo Italy, 110–119. doi:10.1145/345513.345271
@@
- Watson et al. (2019) Katie Watson, Samuel S. Sohn, Sasha Schriber, Markus Gross, Carlos Manuel Muniz, and Mubbasir Kapadia. 2019. StoryPrint: An Interactive Visualization of Stories. In Proceedings of the 24th International Conference on Intelligent User Interfaces (IUI ’19). Association for Computing Machinery, New York, NY, USA, 303–311. doi:10.1145/3301275.3302302
- Wilkinson et al. (2005) Leland Wilkinson, D. Wills, D. Rope, A. Norton, and R. Dubbs. 2005. The Grammar of Graphics (2nd edition ed.). Springer, New York.
- Williams (2018) Eric R. Williams. 2018. How to View and Appreciate Great Movies.
- Yan et al. (2023) Zihan Yan, Chunxu Yang, Qihao Liang, and Xiang ’Anthony’ Chen. 2023. XCreation: A Graph-based Crossmodal Generative Creativity Support Tool. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco CA USA, 1–15. doi:10.1145/3586183.3606826
@@
- Yuan et al. (2022) Ann Yuan, Andy Coenen, Emily Reif, and Daphne Ippolito. 2022. Wordcraft: Story Writing With Large Language Models. In Proceedings of the 27th International Conference on Intelligent User Interfaces (IUI ’22). Association for Computing Machinery, New York, NY, USA, 841–852. doi:10.1145/3490099.3511105
- Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Richmond Y. Wong, Bjoern Hartmann, and Qian Yang. 2023. Why Johnny Can’t Prompt: How Non-AI Experts Try (and Fail) to Design LLM Prompts. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–21. doi:10.1145/3544548.3581388
@@
- Zhang et al. (2023) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023. VISAR: A Human-AI Argumentative Writing Assistant with Visual Programming and Rapid Draft Prototyping. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. ACM, San Francisco CA USA, 1–30. doi:10.1145/3586183.3606800
- Zhao et al. (2025) Zixin Zhao, Damien Masson, Young-Ho Kim, Gerald Penn, and Fanny Chevalier. 2025. Making the Write Connections: Linking Writing Support Tools with Writer’s Needs. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (Yokohama, Kanagawa, Japan) (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 78, 21 pages. doi:10.1145/3706598.3713161
@@
- Zhu (2013) Xiaojin Zhu. 2013. Persistent Homology: An Introduction and a New Text Representation for Natural Language Processing. In Proceedings of the Twenty-Third International Joint Conference on Artificial Intelligence. IJCAI, Beijing, China, 1953–1959.